OpenStudy (anonymous):

Divide these polynomials. (3x3 + x2 - 2x) (x2 + x) A. 3x - 2x B. 3x2 - 3x C. 3x2 - x D. 3x - 2

OpenStudy (anonymous):

@dinamix

OpenStudy (anonymous):

@KendrickLamar2014 hello

OpenStudy (kendricklamar2014):

\[\frac{ 3x3+x2−2x }{ x2+x } = \frac{ 3x2+x−2 }{ x+1}\]

OpenStudy (kendricklamar2014):

\[= \frac{ (3x−2)(x+1) }{ x+1 }\]

OpenStudy (kendricklamar2014):

What does that equal? ^ @2kende

OpenStudy (anonymous):

i dont understand that

OpenStudy (kendricklamar2014):

Lets divide by parts, its really simple: \[\frac{ (3x-2)(x+1) }{ x+1 } = \frac{ (x+1) }{ x+1 } = ?\]

OpenStudy (kendricklamar2014):

That cancels out ^ right?

OpenStudy (anonymous):

what is all this /{[]/) stuff

OpenStudy (kendricklamar2014):

Latex, can you see the numbers?

OpenStudy (anonymous):

yes

OpenStudy (kendricklamar2014):

Do you agree with me that (x+1)/x+1 cancels out?

OpenStudy (anonymous):

i got 2.1 when u put it in the calculator

OpenStudy (kendricklamar2014):

We dont need a calculator for this, only paper and a pencil/pen.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

can we take it little by little

OpenStudy (anonymous):

how do i solve (x+1)/x+1

OpenStudy (anonymous):

does x+1 = x1?

OpenStudy (kendricklamar2014):

\[\frac{ (x+1) }{ x+1 } = 1\]

OpenStudy (kendricklamar2014):

Since your dividing the same number, the answer would be 1

OpenStudy (anonymous):

ohhhhhhhhhhhhh okk

OpenStudy (anonymous):

how would i do it with 2 different numbers?

OpenStudy (kendricklamar2014):

\[\frac{ 3x-2 }{ 1 } = ?\]

OpenStudy (kendricklamar2014):

What would that equal ^

OpenStudy (anonymous):

and where did u get (x+1)/x+1 from

OpenStudy (anonymous):

can we stick wit the first step please until i understand it

OpenStudy (kendricklamar2014):

Sure, the first step was divide: \[\frac{ 3x^3+x^2−2x }{ x^2+x }\]

OpenStudy (kendricklamar2014):

OS is lagging for me, let me refresh the page.

OpenStudy (anonymous):

i have to divide all that? how do i do that?

OpenStudy (kendricklamar2014):

\[\frac{ x^2 }{ x^2 } = \frac{ x }{ x }\]

OpenStudy (anonymous):

can u break this down , \[\frac{ 3x^3 }{ x^2 } = \frac{ 3x^2 }{ x}\] this is the part that is confusing me

OpenStudy (kendricklamar2014):

Ok so, x^3/x^2 all that changes are the exponents.

OpenStudy (anonymous):

why is the exponent 1 less?

OpenStudy (kendricklamar2014):

We are diving all these numbers: \[\frac{ 3x^3+4x^2−4x }{ x^2+2x }\]

OpenStudy (anonymous):

i wanna put all that in the calculator but u said we dont use one

OpenStudy (kendricklamar2014):

No we just have to solve it out

OpenStudy (anonymous):

thats the problem i dont know how to do that

OpenStudy (anonymous):

answer D?

OpenStudy (kendricklamar2014):

I dont have much time, but here is how you solve this problem. All you have to do is divide: \[\frac{ (3x^3+x^2-2x) }{ (x^2+x) }\] \[= \frac{ (3x^2 + 4x-4) }{ x+1 }\] \[\frac{ (3x−2)(x+1) }{ x+1 }\] \[= 3x-2\]

OpenStudy (kendricklamar2014):

Yes, the answer is D